OpenStudy (anonymous):

the product of 2consective room numbers is 342. Find the room numbers. the number of the 1st room, or smaller number is? the number of the 2nd room is?

OpenStudy (anonymous):

You get this equation from the text: n(n+1)=342 which is equivalent to this quadratic equation: n^2+n-342=0

OpenStudy (anonymous):

You know how to solve that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

18,-19

OpenStudy (anonymous):

Right, -19 doesn't make sense here, so n is 18. The consecutive number is then 19. So to check that is correct: 18*19 = 342

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Hope it made sense :D

OpenStudy (anonymous):

it did